# EarthquakeNPP: Benchmark Datasets for Earthquake Forecasting with Neural Point Processes

### v1

@@ Abstract
Classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes for decades. Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over classical models. However, the currently-used benchmark dataset for NPPs does not represent an up-to-date challenge in the seismological community since it lacks a key earthquake sequence from the region and improperly splits training and testing data. Furthermore, initial earthquake forecast benchmarking lacks a comparison to state-of-the-art earthquake forecasting models typically used by the seismological community. To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by a credible implementation of the ETAS model. The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation. In a benchmarking experiment, we compare three spatio-temporal NPPs against ETAS and find that none outperform ETAS in either spatial or temporal log-likelihood. These results indicate that current NPP implementations are not yet suitable for practical earthquake forecasting. However, EarthquakeNPP will serve as a platform for collaboration between the seismology and machine learning communities with the goal of improving earthquake predictability.
@@ 1 Introduction
Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes. While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model (Ogata, 1988; 1998), despite significant growth in available data (Takanami et al., 2003; Shelly, 2017; Ross et al., 2019; White et al., 2019; Mousavi et al., 2020; Tan et al., 2021; Mousavi & Beroza, 2023).
In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility (Du et al., 2016; Omi et al., 2019a; Shchur et al., 2019; Jia & Benson, 2019; Chen et al., 2021; Zhou et al., 2022; Zhou & Yu, 2024). While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community. The benchmark lacks a key earthquake sequence from the region, fails to recreate an operational setting with proper train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
Here, we introduce EarthquakeNPP: a curated collection of datasets designed for benchmarking NPP models in earthquake forecasting, accompanied by a state-of-the-art benchmark model. These datasets are derived from publicly available raw data, which we process and configure within our platform to facilitate meaningful forecasting experiments relevant to stakeholders in the seismology community. Covering various regions of California, these datasets represent typical forecasting zones and encompass data commonly utilized by forecast issuers. Moreover, employing modern techniques, some datasets include smaller magnitude earthquakes, exploring the potential of numerous small events to enhance forecasting performance through flexible NPPs. To unify efforts, we present an operational-level implementation of the ETAS model alongside the datasets, serving as a benchmark for NPPs.
Although initial benchmarking finds that none of the 3 tested NPP implementations outperform ETAS, EarthquakeNPP aims to serve as a platform for future NPP development. The platform facilitates the generative evaluation procedure used for rigorous benchmarking in the seismology community, directing the impact of future NPPs to stakeholders in seismology. Access to the dataset collection, along with comprehensive documentation and notebooks, can be found at https://anonymous.4open.science/r/EarthquakeNPP-2D51.
@@ 1.1 Related Work
Existing Benchmark Dataset. Chen et al. (2021) introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S. Geological Survey, focusing on Japan from 1990 to 2020. They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset. They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake (Mori et al., 2011), the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0. This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact. Additionally, these events are of significant scientific interest because they provide valuable insights into the earthquake rupture process.
@@
The dataset segments are divided for training, testing, and validation. Instead of a chronological partitioning that mirrors operational forecasting, the segments are assigned in an alternating pattern. This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering (Freed, 2005). Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards it time.
@@
Although earthquakes with magnitudes above 2.5 are considered by Chen et al. (2021), following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset. For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets. Section A.2 describes the biases incurred from such data missingness.
@@
Chen et al. (2021) benchmark their model against another spatio-temporal model, Neural Jump SDEs (Jia & Benson, 2019), and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark. Subsequent papers adopting this benchmark (Zhou et al., 2022; Yuan et al., 2023; Zhou & Yu, 2024) similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
Temporal-NPP Benchmarking on Earthquake Data. Two existing works benchmark NPPs for earthquake forecasting within the seismology community. The first by Dascher-Cousineau et al. (2023) extends a temporal-only NPP from Shchur et al. (2019) to include earthquake magnitudes. The second by Stockman et al. (2023) extends another temporal-only model by Omi et al. (2019a) to target larger magnitude events. Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline. Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
@@
Benchmarking within the Seismology Community. Model comparison has been crucial in the development of earthquake forecasting models since their inception (Kagan & Knopoff, 1987; Ogata, 1988). The Collaboratory for the Study of Earthquake Predictability (CSEP) (Michael & Werner, 2018; Schorlemmer et al., 2018; Savran et al., 2022; Iturrieta et al., 2024) (https://cseptesting.org/ ) aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally. CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day). These simulated forecasts are compared by discretizing time and space intervals, with test statistics calculated for event counts, magnitudes, locations, and times. The simulation-based approach allows the inclusion of generative models that don’t output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
1.2 Scope of this work
Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development (Ogata, 1988; Harte, 2015). Our platform uses this metric in the NPP benchmarking experiment and provides detailed guidance on CSEP’s simulation-based procedure, enabling future NPP implementations and evaluations within CSEP experiments.
The goal of this work is to allow Machine Learning researchers to have seismological impact by defining a baseline target for which NPP models can be compared to state-of-the-art domain-based models. NPPs that can generate log-likelihoods are in scope, whilst those that do not (e.g. Yuan et al., 2023; Li et al., 2023) are out of scope because a valid score does not currently exist. The popular next-event point prediction metrics (e.g. Root Mean Square Error (RMSE) and related scores) are considered to be flawed and misleading for seismological prediction (Hodson, 2022), because the predictive distribution is strongly skewed and therefore far from Gaussian. To have seismological relevance, authors of NPP models are challenged to implement long-term predictions using CSEP’s evaluation procedure, benchmarking against the reported performance for the ETAS model.
@@ 2.1 Spatio-Temporal Point Processes
A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\in\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$. This process is typically defined by a non-negative conditional intensity function
$\lambda(t,\mathbf{x}|\mathcal{H}_{t}):=\lim_{\Delta t,\Delta\mathbf{x}\rightarrow 0}\frac{\mathbb{E}\left[N([t,t+\Delta t)\times B(\mathbf{x},\Delta\mathbf{x})|\mathcal{H}_{t}\right]}{|B(\mathbf{x},\Delta\mathbf{x})|},$ (1)
@@
Most models specify the conditional intensity function, though some (e.g. Shchur et al., 2019; Chen et al., 2021; Yuan et al., 2023), directly model this pdf. Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $[T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
[…]
For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric. This is consistent with a wealth of literature in the seismology community (see Zechar et al., 2010, and references therein) as well as the wider general point process literature (Daley & Vere-Jones, 2004), which now more recently includes neural point processes (Shchur et al., 2021). The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident. Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $[T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function. Comparing the mean log-likelihood per event provides the information gain from one model to another (Daley & Vere-Jones, 2004).
2.2 ETAS
The Epidemic Type Aftershock Sequence (ETAS) model (Ogata, 1998) is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space. It has been adopted for operational earthquake forecasting by government agencies in California (Milner et al., 2020), New-Zealand (Christophersen et al., 2017), Italy (Spassiani et al., 2023), Japan (Omi et al., 2019b) and Switzerland (Mizrahi et al., 2024), and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments (e.g. Woessner et al., 2011; Rhoades et al., 2018; Taroni et al., 2018; Cattania et al., 2018; Mancini et al., 2019; 2020; 2022). The general formulation of the model is
$\lambda(t,\mathbf{x}|\mathcal{H}_{t};\mathbf{\theta})=\mu+\sum_{i:t_{i}<t}g(t-t_{i},||\mathbf{x}-\mathbf{x}_{i}||^{2}_{2},m_{i}),$ (4)
where $\mu$ is a constant background rate of events, $g(\cdot,\cdot,\cdot)$ is a non-negative excitation kernel which describes how past events contribute to the likelihood of future events and $m_{i}$ are the associated magnitudes of each event. The equivalent formulation as a Hawkes branching process accompanies a causal branching structure $\mathbf{B}$. This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g. via dynamic wave triggering (Brodsky & van der Elst, 2014) and static stress triggering (Gomberg, 2018; Mancini et al., 2020).
[…]
@@ 2.2 ETAS
This avoids the need to numerically approximate the integral term in the likelihood, provides more stability during estimation and simultaneously estimates the causal structure.
The formulation of the ETAS model we present with the EarthquakeNPP datasets is implemented in the etas python package by Mizrahi et al. (2022). It defines the triggering kernel as
$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{cut})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{cut})}\right)^{1+\rho}},$ (6)
where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$.
@@ 3 EarthquakeNPP Datasets
The EarthquakeNPP datasets encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021. California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms (Gerstenberger et al., 2004; Field, 2007; Field et al., 2021). It encompasses the San Andreas fault plate boundary system (Zoback et al., 1987) and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.

### v2

# EarthquakeNPP: A Benchmark for Earthquake Forecasting with Neural Point Processes

@@ Abstract
For decades, classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes. Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over such classical models. However, the currently-used benchmark for NPPs does not represent an up-to-date challenge in the seismological community, since it contains data leakage and omits the largest earthquake sequence from the region. Additionally, initial earthquake forecasting benchmarks fail to compare NPPs with state-of-the-art forecasting models commonly used in seismology. To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by an implementation of the state-of-the-art forecasting model: ETAS. The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation. Benchmarking experiments, using both log-likelihood and generative evaluation metrics widely recognised in seismology, show that none of the five NPPs tested outperform ETAS. These findings suggest that current NPP implementations are not yet suitable for practical earthquake forecasting. Nonetheless, EarthquakeNPP provides a platform to foster future collaboration between the seismology and machine learning communities.
@@ 1 Introduction
Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes. While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model [53, 54], despite significant growth in available data [79, 72, 62, 90, 51, 80, 50].
In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility [9, 56, 69, 31, 5, 97, 96]. While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community. The benchmark omits the largest earthquake sequence from the region, introduces data leakage with non-sequential train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
Here, we introduce EarthquakeNPP: a curated collection of datasets designed for benchmarking NPP models in earthquake forecasting, accompanied by a state-of-the-art benchmark model. These datasets are derived from publicly available raw data, which we process and configure within our platform to facilitate meaningful forecasting experiments relevant to stakeholders in the seismology community. Covering various regions of California, these datasets represent typical forecasting zones and encompass data commonly utilized by forecast issuers. Moreover, employing modern techniques, some datasets include smaller magnitude earthquakes, exploring the potential of numerous small events to enhance forecasting performance through flexible NPPs. To unify efforts, we present an operational-level implementation of the ETAS model alongside the datasets, serving as the benchmark for NPPs.
Although initial benchmarking shows that none of the five tested NPP implementations outperform ETAS, EarthquakeNPP is designed to support ongoing model development and evaluation. In addition to the standard log-likelihood metric common in the NPP literature, the platform incorporates the generative evaluation procedures used in seismology for more rigorous benchmarking. This ensures that future NPPs (and other models such as time series approaches [88] and Bayesian point processes [68]) can have direct relevance to seismological stakeholders. All datasets, experiments, and documentation are available at https://github.com/ss15859/EarthquakeNPP/tree/main.
@@ 1.1 Related Work
Benchmarking by the NPP Community. Chen et al. [5] introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S. Geological Survey, focusing on Japan from 1990 to 2020. They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset. They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake [49], the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0. This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact.
@@
The dataset segments are divided for training, testing, and validation. Instead of a chronological partitioning that mirrors operational forecasting, the segments are assigned in an alternating pattern. This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering [13] - a form of data leakage. Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards in time.
@@
Although earthquakes with magnitudes above 2.5 are considered by Chen et al. [5], following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset. For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets. Section A.2 describes the biases incurred from such data missingness.
@@
Chen et al. [5] benchmark their model against another spatio-temporal model, Neural Jump SDEs [31], and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark. Subsequent papers adopting this benchmark [97, 94, 96] similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
@@
Benchmarking by the Seismology Community. Model comparison has been crucial in the development of earthquake forecasting models since their inception [34, 53]. The Collaboratory for the Study of Earthquake Predictability (CSEP) [41, 66, 64, 30] [https://cseptesting.org/ ] aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally. CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day). These simulated forecasts are compared by discretizing time and space intervals, with test statistics calculated for event counts, magnitudes, locations, and times. The simulation-based approach allows the inclusion of generative models that don’t output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
Two existing works benchmark NPPs for earthquake forecasting within the seismology community. The first by Dascher-Cousineau et al. [8] extends a temporal-only NPP from Shchur et al. [69] to include earthquake magnitudes. The second by Stockman et al. [77] extends another temporal-only model by Omi et al. [56] to target larger magnitude events. Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline. Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
@@ 2.1 Spatio-Temporal Point Processes
A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\subseteq\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$. This process is typically defined by a non-negative conditional intensity function
$\lambda(t,\mathbf{x}|\mathcal{H}_{t}):=\lim_{\Delta t,\Delta\mathbf{x}\rightarrow 0}\frac{\mathbb{E}\left[N([t,t+\Delta t)\times B(\mathbf{x},\Delta\mathbf{x})|\mathcal{H}_{t}\right]}{\Delta t|B(\mathbf{x},\Delta\mathbf{x})|},$ (1)
@@
$p(t,\mathbf{x}|\mathcal{H}_{t_{i}})=\lambda(t,\mathbf{x}|\mathcal{H}_{t_{i}})\cdot\exp\left(-\int_{t_{i}}^{t}\int_{\mathcal{S}}\lambda(s,\mathbf{z}|\mathcal{H}_{s})d\mathbf{z}ds\right).$ (2)
Most models specify the conditional intensity function, though some [e.g. 69, 5, 94] directly model this pdf. Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
$\log p(\mathcal{H}_{T})=\underbrace{\sum_{i=0}^{n}\log\lambda(t_{i}|\mathcal{H}_{t_{i}})-\int_{T_{0}}^{T_{1}}\int_{\mathcal{S}}\lambda(s,\mathbf{z}|\mathcal{H}_{s})d\mathbf{z}ds}_{\text{Temporal log-likelihood}}+\underbrace{\sum_{i=0}^{n}\log f(\mathbf{x}_{i}|t_{i},\mathcal{H}_{t_{i}})}_{\text{Spatial log-likelihood}},$ (3)
For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric. This is consistent with a wealth of literature in the seismology community [see 95, and references therein] as well as the wider general point process literature [7], which now includes neural point processes [70]. The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident. Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function. Comparing the difference in mean log-likelihood per event provides the information gain from one model to another [7].
Point processes are the dominant modeling approach in the seismology community, used extensively in both real-time operational earthquake forecasting [47] and established benchmarking experiments (CSEP) [81, 60]. The point process representation of earthquake data aligns naturally with their occurrence as discrete events in time [33]. Furthermore, this modeling approach is favored over discretized forecasting models (e.g., time series) because it eliminates the need for optimizing binning strategies and allows for immediate updates, rather than waiting until the end of a time bin — a delay that could miss critical, potentially damaging events.
2.2 ETAS
The Epidemic Type Aftershock Sequence (ETAS) model [54] is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space. It has been adopted for operational earthquake forecasting by government agencies in California [44], New-Zealand [6], Italy [74], Japan [57] and Switzerland [48], and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments [e.g. 93, 60, 81, 4, 38, 39, 40]. The general formulation of the model is
@@ 2.2 ETAS
where $\mu$ is a constant background rate of events, $g(\cdot,\cdot,\cdot)$ is a non-negative excitation kernel which describes how past events contribute to the likelihood of future events and $m_{i}$ are the associated magnitudes of each event. The equivalent formulation as a Hawkes branching process accompanies a causal branching structure $\mathbf{B}$. This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g. via dynamic wave triggering [3] and static stress triggering [16, 39].
Although ETAS can be fit by maximizing the log-likelihood function directly, parameter estimation is typically performed by simultaneously estimating the branching structure $\mathbf{B}$. Veen and Schoenberg [87] developed an Expectation Maximisation (EM) procedure, which maximises the marginal likelihood over the unobserved branching structure, $\log\int p(\mathcal{H}_{T_{1}}|\mathbf{B},\mathbf{\theta})p(\mathbf{B}|\mathbf{\theta})d\mathbf{B}$ through the iteration
@@
This avoids the need to numerically approximate the integral term in the likelihood, provides more stability during estimation, and simultaneously distinguishes background events from triggered events.
Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP. Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans. In red is a fault map from the GEM Global Active Faults Database [78].
The formulation of the ETAS model we present in the EarthquakeNPP benchmark is implemented in the etas python package by Mizrahi et al. [46]. It defines the triggering kernel as
$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{c})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{c})}\right)^{1+\rho}},$ (6)
where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$. This triggering kernel is derived from statistical distributions found through decades of observational studies [84, 83, 85] and several of the learnable parameters have been linked to physical properties of the earthquake rupture process [85, 29].
@@ 3 EarthquakeNPP Datasets
The EarthquakeNPP datasets (Figure 1) encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021. California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms [14, 11, 12]. It encompasses the San Andreas fault plate boundary system [99] and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.